OpenStudy (anonymous):

1.Solve S = 4v2 for v

mathslover (mathslover):

\[\huge{S=4v^2}\] \[\huge{v^2=\frac{S}{4}}\] \[\huge{v=\sqrt{\frac{S}{4}}}\] can u simplify it further ?

mathslover (mathslover):

and also @ayeegabbie WELCOME TO OPENSTUDY . ENJOY ASKING AND ANSWERING QUESTIONS HERE .

OpenStudy (lgbasallote):

replying* not answering :p lol